OpenStudy (anonymous):

the cost of typing a document is directly proportional to the number of pages in the document. If it costs $325 to type a document with 130 pages, how much would it cost to type a document with 85 pages ?

OpenStudy (apoorvk):

HINT: cost per page is constant.

OpenStudy (anonymous):

wait....soo i put it like this? 325/130 = x/85? or do it the k=xy way?

OpenStudy (apoorvk):

exactly!!!! you're right!!

OpenStudy (apoorvk):

the first one i mean.

OpenStudy (anonymous):

oh.....okay then I got $212.5....right?

OpenStudy (apoorvk):

Yup!